OpenStudy (anonymous):

How would you write that the domain includes every real number above and including -1 but not including 6? either set or interval will work Thanks :) :)

OpenStudy (hoblos):

we have the union of two sets here, between -1 and 6, and above 6 let's start with the first part the set of numbers between two numbers a and b including a but excluding b is [a,b[ so what would the set of numbers above and including -1 , and below 6 ?

OpenStudy (anonymous):

[-1,6) or \[-1\le x <6\] right?

OpenStudy (hoblos):

correct for the second part , the set of numbers greater than a number a can be represented by x > a or (a,∞) or ]a,∞[

OpenStudy (hoblos):

so how can we represent numbers above 6 ?

OpenStudy (anonymous):

x > 6 or \[(6,\infty)\]

OpenStudy (hoblos):

correct to join to sets we use U so we get [-1,6) U (6,∞)

OpenStudy (anonymous):

and then can you write it like: Domain =\[\left\{ x \in \mathbb{R}: [-1,6) U (6,\infty) \right\}\]

OpenStudy (anonymous):

Thank You!